# Three-dimensional Imaging of Pion using Lattice QCD: Generalized Parton Distributions

### v1

@@ II.3 Lattice results on the Lorentz-invariant amplitudes
The amplitudes obtained from these two datasets with $P_{z}=0.968$ GeV are shown in Fig. 1. The left two panels show the results obtained using both 2-state and 3-state fits within the Breit and non-Breit frames, respectively. The 2-state fit results are represented by the open symbols, while the 3-state fit results are denoted as the filled symbols, labeled as “2s” and “3s” in the figure, respectively. One can see that the contamination from the excited states is apparent in the Breit frame, so the 3-state fit is needed. For the non-Breit frame, it can be seen that these two kinds of fit results are comparable within 1-$\sigma$ error, and the 2-state fit results are more stable. In both frames, we find that $A_{3}$ is zero within error, as expected. In the right panel of Fig. 1, we investigate the frame independence of $A_{1}$ by selecting the 3-state fit results for the Breit frame and the 2-state fit results for the non-Breit frame. Furthermore, since we observe that $A_{3}$ is compatible with zero within errors also for the non-Breit frame, we perform the extraction of $A_{1}$ from the non-Breit frame based on setting $A_{3}=0$ and using the $\gamma_{t}$ component. The results from this analysis are shown as the green triangles in the right panel of Fig. 1. As depicted in the figure, the non-Breit frame results, both with and without setting $A_{3}=0$, exhibit good agreement and also are consistent with the Breit frame results.
@@
II.4 Renormalization
As we mentioned before, the renormalization of the qGPD matrix elements is crucial for removing the UV divergences from the Wilson line as well as for matching to the light-cone GPDs, which are usually defined in the $\overline{\rm MS}$ scheme. Since the non-local quark bilinear operator is multiplicatively renormalizable, if we call the matrix elements extracted directly from the lattice calculations as the bare matrix elements, the relation between the bare and renormalized matrix elements can be expressed as [75, 76, 77, 49]
$M^{\rm B}(z,a)=Z(a)e^{-\delta m(a)|z|}e^{-\bar{m}_{0}|z|}M^{\rm R}(z,a),$ (14)
where the superscripts “B” and “R” denote the “bare” and “renormalized” quantities, respectively, $Z(a)$ includes the $z$-independent logarithmic divergence, the term with $\delta m(a)$ accounts for the linear divergence, and the term with $\bar{m}_{0}$ is introduced to address the scheme dependence of $\delta m$ and to match the lattice scheme to the $\overline{\rm MS}$ scheme [78, 70, 57, 58]. Theoretically, $\bar{m}_{0}$ should be a constant and independent of $z$.
The hybrid scheme renormalization is defined by merging the ratio scheme for short distances with the explicit subtraction of self-energy divergences in the Wilson line for long distances [49]
@@ II.4 Renormalization
At short distances, all the divergences can be directly canceled by such a ratio. However, for long distances, in order to perform the renormalization, we first need to determine the values of $\delta m$ and $\bar{m}_{0}$. In line with our previous studies, we determine $\delta m$ using lattice QCD results on the static quark-antiquark potential and the free energy of a static quark at non-zero temperatures [79]: $a\delta m=0.1508(12)$ for $a=0.04$ fm lattice [70]. The value of $\bar{m}_{0}$ can be obtained using the bare matrix element results at zero momentum and zero momentum transfer. Namely, by comparing the lattice computations at $P_{z}=t=0$ with their corresponding $\overline{\rm MS}$ Operator Product Expansion (OPE) expressions for such a ratio
@@
$e^{(\delta m+\bar{m}_{0})\delta z}\frac{M^{\rm B}(z+\delta z)}{M^{\rm B}(z)}=\frac{C_{0}(\alpha_{s}(\mu_{0}(z+\delta z)),\mu_{0}^{2}(z+\delta z)^{2})}{C_{0}(\alpha_{s}(\mu_{0}(z)),\mu_{0}^{2}z^{2})}\times\exp\left[\int_{\alpha_{s}(\mu_{0}(z+\delta z))}^{\alpha_{s}(\mu_{0}(z))}{d\alpha_{s}(\mu^{\prime})\over\beta[\alpha_{s}(\mu^{\prime})]}\gamma_{O}[\alpha_{s}(\mu^{\prime})]\right],$ (17)
@@
where $\gamma_{O}$ is the anomalous dimension of the quark bilinear operator, known up to three loops [80], and the expressions for $\alpha_{s}$ and $\beta$ can be found in Appendix B, we can extract the values of $\bar{m}_{0}$. Here, we set $|\delta z|/a=1$. The use of the zero momentum matrix element results in only the 0-th Wilson coefficient $C_{0}$ remaining, which is calculated up to the NNLO [47, 54]. Furthermore, we have incorporated RGR [55] at the next-to-next-to-leading-logarithm (NNLL) accuracy. The RGR allows us to effectively sum the large logarithmic terms arising in the perturbative expansions and to evolve the running coupling from the physical scale $\mu_{0}=2\kappa e^{-\gamma_{E}}/|z|$ to the factorization scale $\mu$ = 2 GeV. Here, $\kappa$ is a proportionality constant, and $\gamma_{E}$ is the usual Euler constant. In principle, we should choose $\kappa\sim 1$ to match the natural physical scale $1/|z|$, but we found that Eq. (17) with $\kappa<1$ cannot describe the lattice data with a constant $\bar{m}_{0}$ in $z$. Therefore, to estimate the scale variation uncertainty, we choose our central value of $\kappa$ as 1.414 and vary it between 1 and 2, where the strong coupling constant remains reasonably small ($\alpha_{s}<$ 0.3) at the short distances we considered. Additionally, we include LRR [58, 57] in the Wilson coefficient that captures the dominant contributions from renormalons and improves the convergence properties of the perturbative series [58].
In Fig. 2, we show the results of $\bar{m}_{0}$ obtained with NNLO(NLO)+LRR and NNLO(NLO)+LRR+RGR Wilson coefficients. As mentioned before, $\bar{m}_{0}$ is excepted to be a constant and independent of $z$. However, in practice, this is not always the case for all $z$ values, as illustrated in Fig. 2. For the two smallest values of $z$, we observe lattice artifacts in the determination of $\bar{m}_{0}$, as expected. However, the dependence of $\bar{m}_{0}$ on $z$ becomes much milder for larger $z$ values. The only exception is the case $\kappa=1$, where the running coupling becomes very large for $z>$ 0.2 fm, causing the perturbative expansions for the Wilson coefficients to break down. The values of $\bar{m}_{0}$ determined using NLO+LRR also show some dependence on $z$ due to missing higher-order corrections. However, thanks to LRR, the dependence of $\bar{m}_{0}$ on $z$ is significantly reduced already at this level compared to our prior analyses of the pion PDFs [70]. All other determinations of $\bar{m}_{0}$ obtained with $\kappa\geq 1.414$ are consistent with each other and show very small $z$-dependence for $z>0.08$ fm. In the following, we will take the $\bar{m}_{0}$ results at $z=0.12$ fm to perform the hybrid-scheme renormalization. For example, $\bar{m}_{0}=0.1922(3)$ is taken in the case of NNLO+LRR+RGR ($\kappa=1$). We also choose $z_{s}=0.12$ fm as our default choice for the hybrid scheme.
@@
Figure 2: The additional Wilson line renormalization mass parameter $\bar{m}_{0}$, obtained using NNLO(NLO)+LRR and NNLO(NLO)+LRR+RGR expressions for $C_{0}$ at different values of $z$. The filled symbols correspond to NNLO results for $C_{0}$, while the open symbols correspond to NLO results for $C_{0}$.
II.5 Large $z$ extrapolation and Fourier transform
@@
In this work, our aim is to get the valence light-cone GPDs in momentum ($x$) space. In the isospin symmetric limit, the valence light-cone GPDs, as well as qGPDs, are equal to the isovector GPDs, as discussed in Ref. [68]. To obtain the momentum-space isovector (valence) qGPDs, we need to perform a Fourier transform on $\lambda=zP_{z}$ of the coordinate-space isovector qGPDs
@@
In Fig. 3, we present our results of the renormalized matrix elements, denoted by the red square symbols, as a function of $\lambda$. Due to the large statistical errors at large $z$ and the finite volume effects, we can reliably calculate the renormalized matrix elements only up to a certain value of $\lambda$. This prevents us from performing the Fourier transform directly on $M^{\bar{\rm R}}$. To address this problem, we perform an extrapolation of the renormalized matrix elements from a truncation position $z_{t}$ and replace the lattice results with the extrapolated values for $z\geq z_{t}$. We expect that $M^{\bar{\rm R}}$ should vanish exponentially at large $z$ [49, 70], which means that the extrapolation mainly affect the small-$x$ region of GPDs [70]. Therefore, we fit the lattice results of $M^{\bar{\rm R}}$ at large $z$ using the following ansatz
@@
where $\{A,m,d\}$ are the fit parameters. In practice, we choose 6 data points above a critical position $z_{c}$ where the matrix element results become unreliable, such as showing negativity or lacking a decaying trend, to do the fit. Subsequently, employing this decay model with the fitted parameters, we simulate results starting from the truncation position $z_{t}=z_{c}-3$, which is the midpoint of the fit range, and extend towards long distances where the matrix element approaches zero. We constrain the fit parameters with the priors $\{A,m,d\}>0$ and extrapolate the data to $z/a=64$, which can lead to extrapolated results decaying to zero in most cases. However, in the cases of $-t=[0,0.231]$ GeV2 for $P_{z}=1.937$ GeV, due to the slower decay, we need to impose a stricter constraint $m>0.2$ GeV [70] as a fit prior and extrapolate the matrix elements to a longer distance $z/a=200$. The extrapolated results, indicated by the purple circle symbols in Fig. 3, exhibit desired decay behavior at long distances and start to approach zero as the distances become sufficiently large. Therefore, the extrapolation can correct the artificial behavior of the lattice data at large $\lambda$.
Figure 3: The results of the extrapolated together with the original renormalized matrix elements obtained with NNLO+LRR+RGR ($\kappa=1$) coefficients are shown as a function of $\lambda=zP_{z}$.
To obtain the $x$ dependence of the qGPDs, we need to integrate the extrapolated renormalized matrix elements over all values of $\lambda=zP_{z}$ in the Fourier transform. In Fig. 4, we show the valence pion qGPD results at $P_{z}=1.453$ and 1.937 GeV, which are obtained using NNLO+LRR+RGR ($\kappa=1.414$) renormalization. We present results for six $-t$ values corresponding to each $P_{z}$ value. We see that as $-t$ increases, the distributions start at smaller values and decay slower with increasing $x$. If we compare the data with similar $-t$ sizes between these two panels, we can find that there are considerable differences between $P_{z}=1.453$ GeV and $1.937$ GeV results. We expect that the perturbative matching to light-cone GPDs, discussed in the next section, can correct for a significant portion of these differences.
@@
III Numerical results on the valence pion light-cone GPDs
@@
$H(x,\mu,t)$ $=\int_{-\infty}^{\infty}\frac{{\rm d}k}{|k|}\int_{-\infty}^{\infty}\frac{{\rm d}y}{|y|}{\cal C}_{\rm evo}^{-1}\left(\frac{x}{k},\frac{\mu}{\mu_{0}}\right)\times{\cal C}^{-1}\left(\frac{k}{y},\frac{\mu_{0}}{yP_{z}},|y|\lambda_{s}\right)\widetilde{H}(y,P_{z},t,z_{s},\mu_{0})$ (20)
$\approx\sum_{i,j}\frac{\Delta k}{|k_{i}|}\frac{\Delta y}{|y_{j}|}{\cal C}_{\rm evo}^{-1}\left(\frac{x}{k_{i}},\frac{\mu}{\mu_{0}}\right)\times{\cal C}^{-1}\left(\frac{k_{i}}{y_{j}},\frac{\mu_{0}}{y_{j}P_{z}},|y_{j}|\lambda_{s}\right)\widetilde{H}(y_{j},P_{z},t,z_{s},\mu_{0}),$
where $\lambda_{s}=z_{s}P_{z}$, ${\cal C}^{-1}_{\rm evo}$ and ${\cal C}^{-1}$ are the inverses of the DGLAP evolution kernel and the hybrid-scheme matching kernel, respectively. The detailed expressions for these kernels up to the NNLL accuracy are given in Appendix B.
Since the qGPDs vanish at the positive and negative infinities, the integration range is finite in practice. To obtain numerical results on the light-cone GPDs, we can discretize the integration variables and replace the integrals with sums, as shown in the second line of Eq. (20). This procedure can reproduce the exact numerical integration well [70]. In Fig. 5, we demonstrate both qGPD and light-cone GPD results obtained from NNLO+LRR+RGR ($\kappa=1.414$) perturbative matching with three-loop DGLAP evolution [81, 82]. It shows the results from two datasets with different values of $P_{z}$ but similar values of $-t$. As seen from this figure, there is a significant difference in the $x$ dependence between the qGPDs and light-cone GPDs, especially at lower $P_{z}$. The effects of the matching are most significant at small and large values of $x$. Additionally, the $P_{z}$ dependence of the distributions is significantly reduced after matching, as expected. The slight dip in the solid-filled green band is likely a numerical artifact, which should be smooth judging by the smoothness of results nearby. In this figure, we only show the results for the light-cone GPDs for $x>0.15$ for the reasons that will be explained below.
Since a larger $P_{z}$ is favored for better control of power accuracy, we will focus on the case with the largest momentum $P_{z}=1.937$ GeV in the following discussion. We will explore the dependence of the light-cone GPDs on the matching accuracy, renormalization scale, and physical variables.
@@
III.1 The dependence of the light-cone GPDs on the perturbative accuracy of the matching and the renormalization scale
In this subsection, we examine the sensitivity of the pion light-cone GPDs to the perturbative accuracy in the matching and the renormalization scale, aiming to determine the reliability of our results.
In Fig. 6, we present the light-cone GPD results for the largest momentum $P_{z}=1.937$ GeV, obtained using NLO and NNLO perturbative matching with LRR, both with or without RGR. The line-filled bands represent the results obtained without RGR at $\mu=2$ GeV. The solid-filled bands show the results obtained with RGR, considering the variation of $\kappa$. The darker-solid-filled bands display the statistical errors with $\kappa=1.414$ only, while the lighter-solid-filled bands also include additional systematic errors associated with scale variation from $\kappa=1$ to 2. The scale variation of the results is more significant at small and large $x$. We have selected the data with the smallest and largest momentum transfers for a clearer comparison. The NNLO results exhibit smaller scale variation, particularly for data with larger $-t$ values. Moreover, the results with and without RGR are almost consistent in the intermediate $x$ region. However, they have a significant difference for $x<0.2$. Since at low scales, the strong coupling constant and the logarithms diverge in the case of RGR, causing the perturbation theory to break down. Consequently, our light-cone GPD results obtained with RGR are only reliable for $x>0.2$. The RGR also plays a role in the vicinity of $x=1$. In this region, the threshold logs may also be important [55, 83]. However, threshold resummation is beyond the scope of this work. Finally, we note that the results between the NLO and NNLO display convergence, especially for the case with larger $-t$ value.
@@
If one uses the hybrid scheme to perform the renormalization, the resulting light-cone GPDs may depend on the choice of $z_{s}$. Therefore, it is important to check the dependence of our results on $z_{s}$. We compare the valence light-cone GPD results with three different values of $z_{s}$ for $P_{z}=1.937$ GeV in Fig. 7. As a demonstration, we show our results for $-t=0$ GeV2, obtained using NNLO+LRR+RGR. The lines represent the central values obtained with $\kappa=1.414$, while the bands encompass both statistical and systematic errors. This convention for the error bands is used in the subsequent figures as well. We can see that the final valence light-cone GPD results have little dependence on $z_{s}$. The situations are similar for other values of $-t$. Thus, our findings are not sensitive to the choice of $z_{s}$. Since the value of $\bar{m}_{0}$ at $z/a=3$ shows the smallest dependence on the perturbative order and the renormalization scale, we adopt $z_{s}/a=3$ for the following analyses.
Furthermore, the zero momentum transfer case corresponds to the PDF results, allowing for comparison with previous PDF calculations. Specifically, we compare our findings with the global analysis at NLO accuracy (JAM21) [84] and our previous results using the NNLO matching (without LRR or RGR) within the same lattice framework (BNL-ANL21) [70]. Our results agree well with the global fit results for $0.2\lesssim x\lesssim 0.8$. Regarding the comparison with BNL-ANL21, differences for $x<0.2$ are not surprising, as discussed earlier. Additionally, we observe that the results obtained with LRR and RGR in this work show a slight downward shift compared to previous findings in the intermediate $x$ region, while they slightly exceed previous results for $x\lesssim 0.25$.

### v2

@@ 2.3 Lattice results on the Lorentz-invariant amplitudes
The amplitudes obtained from these two datasets with $P_{z}=0.968$ GeV are shown in figure 1. The left two panels show the results obtained using both 2-state and 3-state fits within the Breit and non-Breit frames, respectively. The 2-state fit results are represented by the open symbols, while the 3-state fit results are denoted as the filled symbols, labeled as “2s” and “3s” in the figure, respectively. One can see that the contamination from the excited states is apparent in the Breit frame, so the 3-state fit is needed. For the non-Breit frame, it can be seen that these two kinds of fit results are comparable within 1-$\sigma$ error, and the 2-state fit results are more stable. In both frames, we find that $A_{3}$ is zero within error, as expected. In the right panel of figure 1, we investigate the frame independence of $A_{1}$ by selecting the 3-state fit results for the Breit frame and the 2-state fit results for the non-Breit frame. Furthermore, since we observe that $A_{3}$ is compatible with zero within errors also for the non-Breit frame, we perform the extraction of $A_{1}$ from the non-Breit frame based on setting $A_{3}=0$ and using the $\gamma_{t}$ component. The results from this analyses are shown as the green triangles in the right panel of figure 1. As depicted in the figure, the non-Breit frame results, both with and without setting $A_{3}=0$, exhibit good agreement and also are consistent with the Breit frame results.
@@
2.4 Renormalization
As we mentioned before, the renormalization of the qGPD matrix elements is crucial for removing the UV divergences from the Wilson line as well as for matching to the light-cone GPDs, which are usually defined in the $\overline{\rm MS}$ scheme. Since the non-local quark bilinear operator is multiplicatively renormalizable, if we call the matrix elements extracted directly from the lattice calculations as the bare matrix elements, the relation between the bare and renormalized matrix elements can be expressed as Ji:2017oey ; Ishikawa:2017faj ; Green:2017xeu ; Ji:2020brr
$M^{\rm B}(z,a)=Z(a)e^{-\delta m(a)|z|}e^{-\bar{m}_{0}|z|}M^{\rm R}(z),$ (14)
where the superscripts “B” and “R” denote the “bare” and “renormalized” quantities, respectively, $Z(a)$ includes the $z$-independent logarithmic divergence, the term with $\delta m(a)$ accounts for the linear divergence, and the term with $\bar{m}_{0}$ is introduced to address the scheme dependence of $\delta m$ and to match the lattice scheme to the $\overline{\rm MS}$ scheme LatticePartonCollaborationLPC:2021xdx ; Gao:2021dbh ; Holligan:2023rex ; Zhang:2023bxs . Theoretically, $\bar{m}_{0}$ should be a constant and independent of $z$.
The hybrid scheme renormalization is defined by merging the ratio scheme for short distances with the explicit subtraction of self-energy divergences in the Wilson line for long distances Ji:2020brr
@@ 2.4 Renormalization
At short distances, all the divergences can be directly canceled by such a ratio. However, for long distances, in order to perform the renormalization, we first need to determine the values of $\delta m$ and $\bar{m}_{0}$. In line with our previous studies, we determine $\delta m$ using lattice QCD results on the static quark-antiquark potential and the free energy of a static quark at non-zero temperatures Bazavov:2018wmo : $a\delta m=0.1508(12)$ for $a=0.04$ fm lattice Gao:2021dbh . The value of $\bar{m}_{0}$ can be obtained using the bare matrix element results at zero momentum and zero momentum transfer. Namely, by comparing the lattice computations at $P_{z}=0$ GeV and $t=0$ GeV2 with their corresponding $\overline{\rm MS}$ Operator Product Expansion (OPE) expressions for such a ratio
$e^{(\delta m+\bar{m}_{0})\delta z}\frac{M^{\rm B}(z+\delta z)}{M^{\rm B}(z)}$ $=\frac{C_{0}(\alpha_{s}(\mu_{0}(z+\delta z)),\mu_{0}^{2}(z+\delta z)^{2})}{C_{0}(\alpha_{s}(\mu_{0}(z)),\mu_{0}^{2}z^{2})}$
@@
×$\qquad\times\exp\left[\int_{\alpha_{s}(\mu_{0}(z+\delta z))}^{\alpha_{s}(\mu_{0}(z))}{d\alpha_{s}(\mu^{\prime})\over\beta[\alpha_{s}(\mu^{\prime})]}\gamma_{O}[\alpha_{s}(\mu^{\prime})]\right],$ (17)
@@
where $\gamma_{O}$ is the anomalous dimension of the quark bilinear operator, known up to three loops Braun:2020ymy , and the expressions for $\alpha_{s}$ and $\beta$ can be found in appendix F, we can extract the values of $\bar{m}_{0}$. Here, we set $|\delta z|/a=1$. The use of the zero momentum matrix element results in only the 0-th Wilson coefficient $C_{0}$ remaining, which is calculated up to the NNLO Li:2020xml ; Chen:2020ody . Furthermore, we have incorporated RGR Gao:2021hxl at the next-to-next-to-leading-logarithm (NNLL) accuracy. The RGR allows us to effectively sum the large logarithmic terms arising in the perturbative expansions and to evolve the running coupling from the physical scale $\mu_{0}=2\kappa e^{-\gamma_{E}}/|z|$ to the factorization scale $\mu$ = 2 GeV. Here, $\kappa$ is a proportionality constant, and $\gamma_{E}$ is the usual Euler constant. In principle, we should choose $\kappa\sim 1$ to match the natural physical scale $1/|z|$, but we found that eq. (2.4) with $\kappa<1$ cannot describe the lattice data with a constant $\bar{m}_{0}$ in $z$. Therefore, to estimate the scale variation uncertainty, we choose our central value of $\kappa$ as 1.414, which is still close to 1, and vary it between 1 and 2, where the strong coupling constant remains reasonably small ($\alpha_{s}<$ 0.3) at short distances we considered. More details about the process for selecting the range of $\kappa$ can be found in appendix B. Additionally, we include LRR Zhang:2023bxs ; Holligan:2023rex in the Wilson coefficient that captures the dominant contributions from renormalons and improves the convergence properties of the perturbative series Zhang:2023bxs .
@@
Figure 2: The additional Wilson line renormalization mass parameter $\bar{m}_{0}$, obtained using NNLO(NLO)+LRR and NNLO(NLO)+LRR+RGR expressions for $C_{0}$ at different values of $z$. The filled symbols correspond to NNLO results for $C_{0}$, while the open symbols correspond to NLO results for $C_{0}$. The horizontal band indicates the range of values of $\bar{m}_{0}$ used in our analysis. The vertical indicates the value of $z$ at which $\bar{m}_{0}$ was determined.
In figure 2, we show the results of $\bar{m}_{0}$ obtained with Wilson coefficients at the accuracies of NNLO(NLO)+LRR and NNLO(NLO)+LRR+RGR. To enhance visual clarity, results with varying scale settings of the same order are plotted with a slight horizontal offset. As mentioned before, $\bar{m}_{0}$ is excepted to be a constant and independent of $z$. However, in practice, this is not always the case for all $z$ values, as illustrated in figure 2. For the two smallest values of $z$, we observe lattice artifacts in the determination of $\bar{m}_{0}$, as expected. In contrast, at $\kappa=1.414$ and $2$ the dependence of $\bar{m}_{0}$ on $z$ becomes much milder for larger $z$ values until $z\sim 0.3$ fm where perturbation theory breaks down. The only exception occurs when $\kappa=1$, where the running coupling already becomes very large for $z>0.2$ fm, as indicated in figure 17 in appendix B, causing the perturbative expansion of the Wilson coefficient to break down. The values of $\bar{m}_{0}$ determined using NLO+LRR also show some dependence on $z$ due to missing higher-order corrections. However, thanks to LRR, the dependence of $\bar{m}_{0}$ on $z$ is significantly reduced already at this level compared to our prior analyses of the pion PDFs Gao:2021dbh . All other determinations of $\bar{m}_{0}$ obtained with $\kappa\geq 1.414$ are consistent with each other and show very small $z$-dependence for $z>0.08$ fm. Since the results at $z=0.12$ fm show better consistency among all the strategies, we select the $\bar{m}_{0}$ results at $z=0.12$ fm, as indicated by the intersection area of the gray bands in figure 2, to carry out the hybrid-scheme renormalization. For example, $\bar{m}_{0}=0.1922(3)$ GeV at NNLO+LRR+RGR accuracy with $\kappa=1$. Actually, the slight variation in the choice of $\bar{m}_{0}$ has minimal impact on the renormalized results, as illustrated in appendix C. We also choose $z_{s}=0.12$ fm as our default choice for the hybrid scheme.
2.5 Large $z$ extrapolation and Fourier transform
Figure 3: The extrapolated results, along with the original renormalized matrix elements for the case of $-t=1.095$ GeV2, obtained with NNLO+LRR+RGR($\kappa=1$) coefficients, are presented as a function of $\lambda=zP_{z}$.
@@ 2.5 Large $z$ extrapolation and Fourier transform
In this work, our aim is to get the valence light-cone GPDs in momentum ($x$) space. In the isospin symmetric limit, the valence light-cone GPDs, as well as qGPDs, are equal to the isovector GPDs, as discussed in ref. Izubuchi:2019lyk . To obtain the momentum-space isovector (valence) qGPDs, we need to perform a Fourier transform on $\lambda=zP_{z}$ of the coordinate-space isovector qGPDs
@@
In figure 3, we take the case of $-t=1.095$ GeV2 as an example to present our results of the renormalized matrix elements, denoted by the red square symbols, as a function of $\lambda$. Due to the large statistical errors at large $z$ and the finite volume effects, we can reliably calculate the renormalized matrix elements only up to a certain value of $\lambda$. This prevents us from performing the Fourier transform directly on $M^{\bar{\rm R}}$. To address this problem, we perform an extrapolation of the renormalized matrix elements from a truncation position $z_{t}$ and replace the lattice results with the extrapolated values for $z\geq z_{t}$. We expect that $M^{\bar{\rm R}}$ should vanish exponentially at large $z$ Ji:2020brr ; Gao:2021dbh , which means that the extrapolation mainly affect the small-$x$ region of GPDs which is beyond the region of LaMET prediction Gao:2021dbh . Therefore, we fit the lattice results of $M^{\bar{\rm R}}$ at large $z$ using the following ansatz
@@
where $\{A,m,d\}$ are the fit parameters. In practice, we choose $N$ data points to the left from a critical position $z_{c}$ where the matrix element results become unreliable, such as showing negativity or lacking a decaying trend, to do the fit. Subsequently, employing the decay model (19) with the fitted parameters, we simulate results starting from a truncation position $z_{t}/a=z_{c}/a-2$, which is the best choice to ensure the extrapolation start from large enough distance, and extend towards long distances where the matrix element approaches zero. We constrain the fit parameters with the priors $\{A,m\}>0$ and extrapolate the data to $z/a=64$, which can lead to extrapolated results decaying to zero in most cases. However, in the cases of $-t=[0,0.231]$ GeV2 for $P_{z}=1.937$ GeV, due to the slower decay, we need to impose a stricter constraint $m>0.2$ GeV Gao:2021dbh as a fit prior and extrapolate the matrix elements to a longer distance $z/a=200$. For a clearer description, we display the extrapolated results by the green circles in figure 3, which are calculated using the fitted parameters obtained with $N$ = 3. A more detailed discussion of the fit range comparison is available in appendix D. As illustrated, the extrapolated results exhibit expected decay behavior and approach zero as the distances become sufficiently large. Therefore, the extrapolation can effectively correct the artificial behavior of the lattice data at large $\lambda$.
@@
To obtain the $x$ dependence of the qGPDs, we need to integrate the extrapolated renormalized matrix elements over all values of $\lambda=zP_{z}$ in the Fourier transform. In figure 4, we show the valence pion qGPD results at $P_{z}=1.453$ and 1.937 GeV, which are obtained using NNLO+LRR+RGR ($\kappa=1.414$) renormalization. We present results for six $-t$ values corresponding to each $P_{z}$ value. We see that as $-t$ increases, the distributions start at smaller values and decay slower with increasing $x$. If we compare the data with similar $-t$ sizes between these two panels, we can find that there are considerable differences between $P_{z}=1.453$ GeV and $1.937$ GeV results. We expect that the perturbative matching to light-cone GPDs, discussed in the next section, can correct for a significant portion of these differences.
3 Numerical results on the valence pion light-cone GPDs
@@
$H(x,\mu,t)$ $=\int_{-\infty}^{\infty}\frac{{\rm d}k}{|k|}\int_{-\infty}^{\infty}\frac{{\rm d}y}{|y|}{\cal C}_{\rm evo}^{-1}\left(\frac{x}{k},\frac{\mu}{\mu_{0}}\right){\cal C}^{-1}\left(\frac{k}{y},\frac{\mu_{0}}{yP_{z}},|y|\lambda_{s}\right)\widetilde{H}(y,P_{z},t,z_{s},\mu_{0})$ (20)
$\approx\sum_{i,j}\frac{\Delta k}{|k_{i}|}\frac{\Delta y}{|y_{j}|}{\cal C}_{\rm evo}^{-1}\left(\frac{x}{k_{i}},\frac{\mu}{\mu_{0}}\right){\cal C}^{-1}\left(\frac{k_{i}}{y_{j}},\frac{\mu_{0}}{y_{j}P_{z}},|y_{j}|\lambda_{s}\right)\widetilde{H}(y_{j},P_{z},t,z_{s},\mu_{0}),$
where $\lambda_{s}=z_{s}P_{z}$, ${\cal C}^{-1}_{\rm evo}$ and ${\cal C}^{-1}$ are the inverses of the DGLAP evolution kernel and the hybrid-scheme matching kernel, respectively. The detailed expressions for these kernels up to the NNLL accuracy are given in appendix F.
@@
Since the qGPDs vanish at the positive and negative infinities, the integration range is finite in practice. To obtain numerical results on the light-cone GPDs, we can discretize the integration variables and replace the integrals with sums, as shown in the second line of eq. (20). This procedure can reproduce the exact numerical integration well Gao:2021dbh . In figure 5, we demonstrate both qGPD and light-cone GPD results obtained from NNLO+LRR+RGR ($\kappa=1.414$) perturbative matching with three-loop DGLAP evolution Curci:1980uw ; Moch:2004pa . It shows the results from two datasets with different values of $P_{z}$ but similar values of $-t$. As seen from this figure, there is a significant difference in the $x$ dependence between the qGPDs and light-cone GPDs, especially at lower $P_{z}$. The effects of the matching are most significant at small and large values of $x$. Additionally, the $P_{z}$ dependence of the distributions is significantly corrected after matching, as expected. Since the strong coupling constant becomes ill-defined at low scales in RGR, causing the perturbation theory to break down, we can only calculate the results down to some minimum value of $x$.
Since a larger $P_{z}$ is favored for better control of power accuracy, and a $P_{z}\to\infty$ extrapolation on a single lattice ensemble cannot be rigorously done without controlling the discretization effects, we will focus on the case with the largest momentum $P_{z}=1.937$ GeV in the following discussion. We will explore the dependence of the light-cone GPDs on the matching accuracy, renormalization scale, and physical variables.
3.1 The dependence of the light-cone GPDs on the perturbative accuracy of the matching and the renormalization scale
In this subsection, we will examine the sensitivity of pion light-cone GPDs to both perturbative matching accuracy and renormalization scale, aiming to determine the reliability of our results.
@@
In figure 6, we present the light-cone GPD results for the largest momentum $P_{z}=1.937$ GeV, obtained using NLO and NNLO perturbative matching with LRR, both with or without RGR. The line-filled bands represent the results obtained without RGR at $\mu=2$ GeV. The solid-filled bands show the results obtained with RGR, considering the variation of $\kappa$. The darker-solid-filled bands display the statistical errors with $\kappa=1.414$ only, while the lighter-solid-filled bands also include additional systematic errors associated with scale variation from $\kappa=1$ to 2. The scale variation of the results is more significant at small and large $x$. We have selected the data with the smallest and largest momentum transfers for a clearer comparison. The NNLO results exhibit smaller scale variation, particularly for data with larger $-t$ values. Moreover, the results with and without RGR are almost identical in the intermediate $x$ region but show significant differences at small $x$, as expected. As discussed above, due to the breakdown of perturbation theory, our light-cone GPD results obtained with RGR are only reliable for $x\gtrsim 0.2$. The significant changes in the vicinity of $x=1$ in RGR can be attributed to two factors. First, the quasi-GPD after inverse matching still has a non-vanishing tail at $x\geq 1$ due to the power corrections, which propagates to the final result after DGLAP evolution. Besides, in this region the threshold logarithms in the matching and evolution kernels are important Gao:2021hxl ; Ji:2023pba ; Ji:2024hit . However, threshold resummation is beyond the scope of this work. Nevertheless, according to ref. Ji:2024hit both RGR and threshold resummations have little impact for $x<0.8$ at $P_{z}=1.937$ GeV, so we can focus on our results in the region $0.2\lesssim x\lesssim 0.8$ where the systematics is under control at current perturbative accuracy. Finally, we note that the results between the NLO and NNLO display convergence, especially for the cases with larger $-t$ value.